OpenStudy (mysticfury):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –4 8 4 6 Line 2 x y –1 1 3 5 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (mysticfury):

@sweetburger @ltrout @JDcap

OpenStudy (mysticfury):

@rebeccaxhawaii

OpenStudy (mysticfury):

i think B btw

OpenStudy (welshfella):

these are linear functions of x what are the slopes of each line

OpenStudy (jdcap):

\[y_1=7-\frac{ 1 }{ 4 }x \\ y_2=x+2 \\ x=4;y=6\]

OpenStudy (welshfella):

slope = (y2-y1)/(x2-x1)

OpenStudy (welshfella):

yes thats correct these will intersect once

OpenStudy (mysticfury):

okay so B is correct then "exactly once" @welshfella @JDcap

OpenStudy (jdcap):

Yup